Hi I need help with this problem
The depth of a pool varies directly with how many gallons of water it will hold. If a 6 foot deep pool holds 10,000 gallons of water, a pool that holds 12,500 gallons of water should be __________inches deep, according to this theory.
Thanks
: Hi I need help with this problem
The depth of a pool varies directly with how many gallons of water it will hold. If a 6 foot deep pool holds 10,000 gallons of water, a pool that holds 12,500 gallons of water should be __________inches deep, according to this theory.
Thanks

Answer by stanbon(18792)

The depth of a pool varies directly with how many gallons of water it will hold.
d = kg
------------
If a 6 foot deep pool holds 10,000 gallons of water
6 = k*10000
k = 6/10000
k = 0.0006
EQUATION for your problem:
d = (0.0006)g
----------------
, a pool that holds 12,500 gallons of water should be __________inches deep, according to this theory.
d = (0.0006)(12,500) = 7.5 ft = 90 inches
-------------------------------
Cheers,
Stan H.